# Decentralized Reliability Estimation for Mixnets

### v1

@@ Abstract
Continuous-time decryption mixnets can anonymously route data packets with end to end latency that can be as low as a second, making them usable for a variety of applications. Such mixnets however lack verifiable reliability properties that ensure the correct processing and delivery of packets, while existing verifiability mechanisms are incompatible with scalable low latency continuous-time mixnets due to imposing overheads measuring in minutes to hours. This work addresses this gap by proposing a scheme that can estimate reliability scores for links and nodes forming a continuous-time mixnet where some form of credentials authorize clients to send traffic. The scores can be computed publicly by all participants from a set of measurement packets that are eventually revealed and act as a random sample of the traffic, without affecting mixnet transmission latency for client packets. Our scheme relies on VRF-based routing, a novel primitive that ensures that legitimate client packets follow the routing policy of the mixnet, as well as randomly generating unforgeable measurement packets. We experimentally validate our construction both in unreliable and adversarial settings, demonstrating its feasibility.
@@ 1 Introduction
Mix networks or mixnets, first introduced by Chaum [4], are a basic technique for anonymously routing packets through multiple intermediaries called mix nodes, such that the input and output packets of nodes are unlinkable due to cryptographic transformations and packet reordering. Mixnets route each packet independently, in contrast to Onion Routing [16] networks such as Tor [12] that establish end-to-end bidirectional circuits carrying all the packets of a communication flow — and providing the initiator with a live connection to each routing intermediary. Continuous-time mixnets [29] function by default as a fire-and-forget protocol, where packet delivery is not guaranteed and the sender lacks straightforward mechanisms to check the mix nodes acting as routing intermediaries for their packets. Reliable packet transmissions can be achieved with end-to-end acknowledgments and retransmissions [9]. Such solution however does not reveal which mix node in the path failed to process and forward a packet, neither provides a mechanism to publicly estimate the reliability of individual mix nodes across all the clients’ traffic — information that is critical for network management that accounts for quality of service.
@@
Mixnets can be designed to provide very high assurances with respect to correct processing by all mix nodes and final message delivery by having: (i) round-based communication, with clients normally sending at most one short message per round; (ii) all communications along the route happen via broadcast (or “bulletin board”), thus being publicly verifiable; and (iii) some form of verifiable shuffle at each mix node that allows it to publicly prove (via broadcast) the correctness of the mixing operation, before the next step of routing can proceed. These requirements, and the resulting latency, severely restrict the uses cases where these mixnets can be practically deployed. Voting applications are a common use case for these solutions [15, 27], given their strict integrity and verifiability requirements and their latency tolerance. More scalable and efficient mixnet designs within this paradigm have been proposed for implementing anonymous micro-blogging applications [22] or private client messaging via dead-drops [32, 23]. Even these faster designs however have a routing latency of minutes, which grows with the number of messages routed in the round, making them unsuitable to support diverse applications.
In this paper we consider scalable continous-time mixnets such as Loopix [29] and Nym [9] that can deliver packets with a latency as low as a sub-second, so that the mixnet is practically usable for a broad range of applications including email, instant messaging, file sharing, broadcast of (blockchain) transactions, and general purpose internet access. Furthermore, packet’s end-to-end routing latency should not be affected as the network scales to handle more clients and packets. This can be achieved by mixnet architectures such as Loopix [29] where mix nodes process packets individually, rather than in round-based batches, and where clients can select the per-node delay of their packets from a latency distribution that is compatible with the latency tolerance of their application [18]. The Nym mixnet, [^1] https://explorer.nymtech.net which is an implementation of Loopix with various extensions, currently delays packets by an average of $50$ms at each of three hops, resulting in $500$ms to $800$ms end-to-end latency when network propagation delays are included. Packet processing times take less than $1$ms and their contribution to total latency is negligible in comparison to mixing and propagation time.
Nym [9] includes several additional components compared to Loopix: (1) it implements anonymous credentials to authorize usage of the mixnet in terms of an allowance of packets given to subscribed clients [30]; (2) it includes both a blockchain for long-term data storage and an ephemeral broadcast layer that function as public bulletin board to register nodes, keep track of node staking and rewarding, coordinate network operations and provide a consistent view on the state of the network to all participants; (3) lastly, Nym relies on an incentive mechanism that rewards nodes taking into account their reliability, such that they miss out on rewards if they go offline or otherwise fail to process and forward packets [10]. This is reinforced by a reputation mechanism based on staking that further incentivizes stakeholders to delegate their stake to reliable nodes, which in turn increases their chances of participation in the network. Nodes found to be unreliable both lose immediate rewards and reduce their possibilities to participate in the network in the next epochs. Consistently underperforming nodes eventually lose stake from delegates that switch to more reliable (and profitable) nodes, diminishing the reputation, future rewards, and chances of participation of the underperformers. Note that newly created nodes may need time to build reputation, and thus nodes are incentivized to maintain long-term identities.
Running the Nym incentive mechanism therefore requires computing a ‘reliability’ or ‘performance’ score per node participating in the mixnet. This must be done in a context where packet transmissions between a pair of parties are one-to-one (instead of via shared broadcast) and thus, if the sender and receiver of a transmission disagree, no third parties can verify whether a packet was never sent, sent but not received, or sent and received but not acknowledged as received. Furthermore, in this setup mix nodes are not able to publicly prove that their outputs are legitimate transformations (shuffles) of their inputs. Thus, if packets are dropped or substituted along routes it is generally not trivial to determine who is deviating from the protocol. Due to the lack of existing decentralized solutions, which has long been an open problem in mixnet research [13], current performance estimations in the Nym mixnet rely on a trusted measurement authority that sends test packets to itself to compute performance estimates, which makes all estimates and resulting rewards vulnerable to a faulty or malicious authority.
This work addresses this open problem by proposing a scheme that can estimate the reliability of all the mixnet links and nodes in a decentralized manner — at a fraction of the per-packet overhead and latency offered by verifiable shuffling techniques, while still relying on broadcast to achieve public verifiability. Assuming that clients obtain (anonymous) credentials [30] that entitle them to send an amount of legitimate traffic through the mixnet, our solution further ensures that clients follow the mixnet’s stated routing policy when privately creating legitimate packets. We design our scheme to be compatible with the Nym network, while keeping it as generic as possible so it can be analyzed in a stand alone manner.
@@
Our design is built using a novel primitive that we call VRF-based routing. A VRF (Verifiable Random Function) [26] is a function that maps any input to a pseudorandom value in a way that is publicly verifiable. Based on VRFs we show how we can tie each legitimate packet to a valid client credential and ensure that its route is randomly chosen according to the mixnet’s publicly known routing policy. VRF-based routing further allows turning a mixnet’s packet encoding scheme (such as the widely deployed Sphinx [7] format used by Loopix and Nym) into a verifiably unbiased lottery for measurement packets that are used for network monitoring purposes rather than for carrying client data. The intuition is that of an undercover “secret shopper” that goes to a store to detect malicious or faulty behavior in employees, a concept that has been proposed earlier to probabilistically catch misbehaviour in a mix cascade setting [13]. Measurement packets are indistinguishable from normal packets while in transit, but they can be revealed and verified at a later stage to facilitate legitimate traffic estimates. Taking advantage of this, we introduce a protocol that periodically produces public estimates for the share of legitimate traffic transmitted and dropped at each mixnet link during a time epoch. We further propose a method to estimate node performance scores from link performance.
@@
We empirically evaluate via simulations the proposed protocols. We show that, in the presence of random node failures and given sufficient measurement samples, our node performance estimation protocol produces accurate estimates. Furthermore we show that malicious nodes that attempt to lower their neighbours’ estimates by selectively dropping packets pay a performance cost equivalent to the cost inflicted on their targets, thus neutralizing “creeping death” attacks [13] where adversaries gain relative reputational advantage over time. The proposed protocols are highly scalable with increases in the client base, as the amount of measurements and related overheads remain constant for a desired estimation accuracy and network size, rather than growing proportionally to the amount of client traffic or the size of packets.
@@
Paper organization. We first introduce the problem statement in Sect. 2 by describing the system and threat models and the desired properties. In Sect. 3 we describe the core elements of our scheme, including a protocol that estimates mixnet link performance and the VRF-based routing primitive that is a key building block of our solution. We show in Sect. 4 how link performance estimates can be further processed to compute performance estimates for mix nodes, where we evaluate the accuracy of the proposed methods via simulations, as well as the overhead of a possible implementation. Finally, we review related work in Sect. 5 and conclude with future avenues of research in Sect. 6.
@@ 2.1 System Model
Decryption mixnet model. We consider a decryption mixnet where senders prepare packets by encrypting them with the keys of selected intermediaries, each of which then removes a layer of encryption when routing the packet. We abstract the mixnet topology as a labelled directed graph $G$. Vertices are labelled as either mix nodes, gateways or clients. Edges in $G$ represent the communication links between these entities. We denote the edge from vertex $i$ to $j$ as $e=(i,j)$. Mix nodes have both incoming and outgoing edges to other mix nodes and gateways. Gateways serve both as entry and exit points with respect to the mixnet and have incoming and outgoing edges to both mix nodes and clients. For each entity $j$ we define as $P(j)$ the set of its immediate predecessors and $S(j)$ the set of its immediate successors in $G$. The set of “forward direction predecessors” denoted by $P^{*}(j)$ is defined to be $\emptyset$ for a client $j$, a set of clients in case $j$ is a gateway, and $P^{*}(j)=P(j)\cup_{i\in P^{*}(j)}P^{*}(i)$ for a mix node $j$. A packet route is initiated by a client that is affiliated with a gateway $g_{0}$ and follows a path in $G$ picking outgoing edges and traversing multiple mix nodes. The last mix node in the path is followed by the recipient’s gateway $g_{1}$, who delivers the packet to its intended destination, which could be another client, a service, or a proxy to the open internet.
Client credentials. Users buy a subscription by performing a payment on a smart contract on the blockchain. This subscription has the form of a compact e-cash wallet [31] (with keys $vk_{\mathsf{long}},sk_{\mathsf{long}}$) that contains a bundle of $N$ unlinkable credentials. Each credential $c$ entitles the client to route up to $S_{c}$ packets through the mixnet.
Network and routing policy. We assume for convenience that the mixnet graph $G$ has a layered topology, i.e., $N=LW$ mix nodes are arranged in $L$ layers, with each layer having $W$ nodes. Mix nodes are selected [^2] If more than $N$ nodes are available, the selection criteria accounts for node reliability and reputation, as is the case in Nym, so that underperforming nodes are excluded. from a pool of at least $N$ eligible nodes and assigned randomly to a layer for a time period (epoch), with the node selection being refreshed every new epoch so that nodes rotate positions and underperforming nodes can be replaced. The set of $W_{G}$ gateways that act as interface between clients and the mixnet is assumed to remain relatively stable over time to minimize the need for clients to switch gateways with the change of epoch, though gateways may be replaced, e.g., if they have low popularity among clients or low performance. Nodes in the first layer of the mixnet only accept packets from eligible gateways and clients only spend credentials with eligible gateways of their choice.
@@
Blockchain and broadcast channel. We assume that all nodes have access to a blockchain setup that enables them to post transactions that are stored and available long term. Additionally, nodes can use a “layer 2” ephemeral broadcast channel that enables one-to-many information dissemination for information that only needs to be available for a limited time. Using this layer will be referred to as “broadcasting.” This component is instantiated in the Nym mixnet by the use of a Cosmos-based blockchain with the ‘Ephemera’ layer 2 solution that serves as a general purpose key-value pair database on the blockchain whose state can be reset every few epochs. The code is available online. [^3] https://github.com/nymtech/nym/tree/develop/ephemera
@@ 2.2 Threat Model
  • The adversary may control multiple clients and a limited number of mix nodes and gateways. Malicious mix nodes can adaptively drop or substitute packets as well as lie when broadcasting output values. Active malicious behaviour by gateways is assumed to be utility maximizing: gateways follow any strategy as long as it does not decrease their utility. We allow any utility function to be used for gateways as long as it is monotonically increasing on a number of indicators — defined precisely in Section 3.2.
  • “Honest” nodes (mix nodes and gateways) are divided in a set of reliable nodes that are consistently online and process all received packets and a set of unreliable nodes which experience downtime, throughput limitations, or other faults that result in packet drops. All honest nodes report true output values in every epoch.
  • A credential issuer that issues credentials as requested but is not trusted with respect to the privacy of clients.
  • A blockchain and ephemeral broadcast channel that cannot be subverted by the adversary, i.e., they satisfy consistency: honest parties have the same view regarding their contents (perhaps subject to a lag between them due to catching up with protocol messages at different times), as well as liveness, i.e., given sufficient time after a transaction is posted either on the ledger or the ephemeral broadcast channel, the transaction is available for all to have access to. Note that in the case of the ephemeral broadcast channel, unreliable honest parties may miss transmitted messages, while in the case of blockchain all honest parties eventually catch up on all posted transactions.
@@ 2.3 Properties
There are two cryptographic algorithms that typify every mixnet: the packet encoding algorithm that prepares content to be routed via the mixnet and a packet processing algorithm that is executed at every hop in the packet route. In a paid service setting, these two algorithms are further equipped with the characteristic that there is a limited amount of bandwidth that a user is authorized to use. Prior to running the packet encoding algorithm, clients obtain their e-cash wallet which provides them with an (anonymous divisible) bundle of $N$ credentials; subsequently, they approach a gateway and, using the divisibility property of their subscription, spend one of the $N$ credentials that will entitle them to submit $S_{c}$ packets to that gateway, who in turn forwards them to the mix nodes. Packets that originate from a particular credential and gateway and are encoded correctly using the packet encoding algorithm are called legitimate packets. Note that a legitimate packet is well-defined given the gateway identity, the credential keys and the random coins used by client and gateway assuming both are running the mixnet encoding algorithm.
Given this feature, for each epoch, we associate the following quantities with a link $e$: ${s}_{e}$ is the number of legitimate packets that were successfully transmitted, i.e., sent and received across the link $e$, and ${d}_{e}$ is the number of legitimate packets dropped in the link. We consider that ${d}_{(c,g)}=0$ in the link $e=(c,g)$ between client and gateway, and that ${s}_{(c,g)}$ thus includes all the packets generated from the client’s credential $c$. At each next step of the path, the quantities ${s}_{e}$ and ${d}_{e}$ of the incoming and outgoing links of a gateway or node $j$ satisfy the following relation.
$\sum_{i\in P(j)}{s}_{(i,j)}=\sum_{k\in S(j)}{s}_{(j,k)}+{d}_{(j,k)}$ (1)
@@
A mixnet with reliability estimation (REst-mixnet) is a mixnet scheme over a graph $G$ where (1) time is divided in epochs; (2) in each epoch a publicly known number of legitimate packets are introduced through each gateway; and (3) given an edge $e=(i,j)$, all participants are able to estimate the link’s performance ${\rho}_{e}$, defined as the fraction of legitimate packets successfully transmitted from $i$ to $j$, relative to all the legitimate packets received by $i$ to be forwarded to $j$; i.e., including both the ${s}_{e}$ packets that were successfully received by $j$ and the ${d}_{e}$ packets that were dropped before being registered by $j$:
@@
An estimation $\hat{\rho}_{e}\approx\rho_{e}$ can be obtained by anyone who reads the broadcast channel. For a given confidence level (Z-score), the accuracy of the estimation is bounded by a per-link maximum error $\epsilon_{e}$ which is computed in tandem with $\hat{\rho}_{e}$.
In other words in a REst mixnet, parties are able to estimate the rate of successfully transmitted legitimate packets in individual links and even along multi-link paths of interest in the mixnet graph. At the same time, we require the following two security properties.
@@
Privacy. Finally, the main objective of any mixnet is to provide unlinkability between input and output packets, assuming an adversary that knows the sender and receiver sets, observes traffic in all the network links (including between clients and gateways) and even controls some of the participants. The reliability estimation mechanism should not leak any additional information to the adversary that could be useful to deanonymize packets.
3 Performance Estimation Protocol
We design a REst mixnet by introducing indistinguishable measurement packets into the traffic of the mixnet that are subsequently opened and traced along the paths in the mixnet graph. We lay out the description of our protocol in two parts. First, we describe a sampling protocol that generates measurement traffic and enables mix nodes to present evidence regarding their performance; this allows the performance estimation algorithm that pulls together the broadcast information to estimate the ratio of successfully transmitted packets in each link and its associated error given a confidence interval; this approach works in a setting where clients and gateways are semi-honest. Second, we introduce VRF-based routing, which is a cryptographic tool that generates unforgeable measurement packets according to the routing policy and limits attacks by malicious clients and gateways.
3.1 Link performance estimation
We start by describing our link estimation protocol in the simpler setting where clients and gateways are (semi) honest. Then, we show how VRF-based routing can mitigate deviations by clients and gateways.

### v2

# Decentralized Reliability Estimation for Low Latency Mixnets

@@ Abstract
While there exist mixnets that can anonymously route large amounts of data packets with end to end latency that can be as low as a second, combining this level of performance with strong verifiability and reliability properties that ensure the correct processing and delivery of packets has proved challenging. Indeed, existing verifiability mechanisms are incompatible with scalable low-latency operation due to imposing significant latency overheads measuring in minutes to hours, hence severely limiting the variety of applications mixnets can serve. We address this important gap by proposing a scheme that can estimate reliability scores for a mixnet’s links and nodes in a decentralized manner with essentially optimal complexity that is independent of the total traffic routed through the mixnet. The scores can be computed publicly by all participants from a set of measurement packets that are eventually revealed and act as a random sample of the traffic, without affecting mixnet transmission latency for client packets or incurring significant bandwidth overhead. Our scheme assumes client credentials and relies on VRF-based routing, a novel primitive that ensures that legitimate client packets follow the routing policy of the mixnet, as well as randomly generating unforgeable measurement packets. We experimentally validate our construction both in unreliable and adversarial settings, demonstrating its feasibility.
@@ 1 Introduction
Mix networks or mixnets, first introduced by Chaum [5], are a basic technique for anonymously routing packets through multiple intermediaries called mix nodes, such that the input and output packets of nodes are unlinkable due to cryptographic transformations and packet reordering. Mixnets route each packet independently, each one paving its own path through the mixnet. This is in sharp contrast to other privacy enhancing techniques such as Onion Routing [20] and networks such as Tor [15] that establish end-to-end bidirectional circuits carrying all the packets of a communication flow — in this way providing the initiator with a live connection to each routing intermediary and the ability to estimate relays’ reliability directly.
Mixnet design follows two major approaches to reorder packets: continuous-time [23, 32, 11], where mix nodes apply random delays to individual packets; and batch-based [5, 4, 25], where mix nodes shuffle batches of $N$ packets. Continuous-time mixnets can deliver packets with low latency but they function by default as a fire-and-forget protocol, where packet delivery is not guaranteed and the sender lacks straightforward mechanisms to check the mix nodes acting as routing intermediaries for their packets. Reliable packet transmissions can be achieved with end-to-end acknowledgments and retransmissions [11]. Such solutions however do not reveal which mix node in the path failed to process and forward a packet, nor provide a mechanism to publicly estimate the reliability of mix nodes across all the clients’ traffic — information that is critical for network management that accounts for quality of service.
@@
Batch-based mixnets [37, 30] on the other hand, can offer a higher level of assurance but at substantial cost in terms of overhead and practical usage constraints. Such mixnets offer reliability typically by having (i) round-based communication, with clients normally sending at most one short message per round; (ii) all communications along the route happen via broadcast (or “bulletin board”), thus being publicly verifiable; and (iii) some form of verifiable shuffle at each mix node that allows it to publicly prove (via broadcast) the correctness of the mixing operation, before the next step of routing can proceed. The overhead to cover these requirements severely restricts the uses cases where these mixnets can be practically deployed. Voting applications are a common use case [19, 30], given their strict integrity and verifiability requirements and their latency tolerance. More efficient mixnet designs within this paradigm have been proposed for implementing anonymous micro-blogging applications [25] or private client messaging via dead-drops within the mixnet for message delivery [34, 26]. Even these faster designs however have a routing latency of minutes, which grows with the number of messages routed in the round, making them unsuitable to scale and support diverse applications.
This state of affairs suggests a critical gap: for mixnets to be practically usable for a broad range of applications including email, instant messaging, file sharing, blockchain transactions, and general purpose internet access, both reliability estimation and sub-second latency are needed — a combination that so far has been elusive and seemingly infeasible. Furthermore, packets’ end-to-end routing latency should not be affected as the network scales to handle more clients and packets and the reliability estimation should be publicly verifiable and decentralized so that no system component becomes a single point of failure in this estimation.
Our contributions. We propose a scheme that estimates the reliability of a mixnet’s links and nodes in a decentralized manner — at a fraction of the per-packet overhead and latency offered by verifiable shuffling techniques, while still relying on broadcast to achieve public verifiability. To maintain low latency we focus on continuous-time mixnets such as Loopix [32] and Nym [11], which deliver packets with a latency as low as a sub-second, [^1] Nym (https://explorer.nymtech.net) is an implementation of Loopix with various extensions, including client credentials [33] and an incentive mechanism for nodes where rewards depend on node reliability [12]. Nym currently delays packets by an average of $50$ms at each of three mixing hops, resulting in $500$ms to $800$ms end-to-end latency when network propagation delays are included. Packet processing takes less than $1$ms and is thus negligible compared to mixing and propagation time. but note that our methods can also be applied to batch-based mixnets as well. We design our scheme to be compatible with the Nym network, while keeping it as generic as possible so it can be analyzed in a stand-alone manner.
@@
Our design is built using a novel primitive that called VRF-based routing. A VRF (Verifiable Random Function) [29] maps any input to a pseudorandom value in a way that is publicly verifiable. Based on VRFs we show how we can tie packets to client credentials and ensure that their route is randomly chosen according to the mixnet’s publicly known routing policy. VRF-based routing allows turning a mixnet’s packet encoding scheme (such as the widely deployed Sphinx [9] format used by Loopix and Nym) into a verifiably unbiased lottery for measurement packets that are used for network monitoring purposes rather than merely carrying data. The intuition is that of an undercover “secret shopper” that goes to a store to detect malicious or faulty behavior in employees, a concept that has been proposed earlier to probabilistically catch misbehaviour in a mix cascade setting [17]. Measurement packets are indistinguishable from normal packets while in transit, but they can be revealed and verified at a later stage to facilitate legitimate traffic estimates. Taking advantage of this, we introduce a protocol that periodically produces public estimates for the share of legitimate traffic transmitted and dropped at each mixnet link during a time epoch. We further propose a method to estimate node reliability scores based on the packets dropped in their incoming and outgoing links.
To understand the power of the approach consider a communication link involved in $n$ packet transmissions, exhibiting some arbitrary reliability pattern of dropping packets over a period. Using VRF-based routing we can disperse a series of indistinguishable measurements within the packet sequence. Given the indistinguishability of the packets, one can think of the link as implementing a lossy channel that acts as a coin flip of unknown bias $\rho$. The opening of the measurement packets results in a series of publicly observable coin flips of this $\rho$-biased coin. From this observed sequence of coin flips, the conditional distribution of $\rho$ is a Beta distribution with variance $O(1/k)$ where $k$ is the number of measurement packets. It follows that we can estimate $\rho$ up to error $\sim 1/\sqrt{k}$ with broadcast communication complexity equal to $O(k)$, i.e., independently of the total amount of traffic routed in the network, and only proportional to the amount of measurements. Furthermore, this is essentially optimal as such number of coin flips is generally needed for error estimation at that level.
@@
We empirically evaluate via simulations the proposed protocols. We show that, in the presence of node failures and given sufficient measurement samples, our node reliability estimation protocol produces accurate estimates. Furthermore we show that malicious nodes that attempt to lower their neighbours’ estimates by selectively dropping packets pay a reliability cost equivalent to the cost inflicted on their targets, thus neutralizing “creeping death” attacks [17] where adversaries gain relative reputational advantage over time. The proposed protocols are highly scalable with increases in the client base, as the amount of measurements and related overheads remain constant for a desired estimation accuracy and number of network links, rather than growing proportionally to the amount of client traffic or the size of packets.
@@
Paper organization. We first introduce the problem statement in Sect. 2 by describing the system and threat models and the desired properties. In Sect. 3 we describe the core elements of our scheme, including a protocol that estimates mixnet link reliability and the VRF-based routing primitive that is a key building block of our solution. We show in Sect. 4 how link reliability estimates can be further processed to compute reliability estimates for mix nodes, where we evaluate the accuracy of the proposed methods via simulations, as well as the overhead of a possible implementation. Finally, we review related work in Sect. 5 and conclude in Sect. 6.
@@ 2.1 System Model
Decryption mixnet model. We consider a continuous-time decryption mixnet where senders prepare packets by encrypting them with the keys of selected intermediaries, each of which then removes a layer of encryption when routing the packet. We abstract the mixnet topology as a labelled directed graph $G$. Vertices are labelled as either mix nodes, gateways or clients. Edges in $G$ represent the communication links between these entities. We denote the edge from vertex $i$ to $j$ as $e=(i,j)$. Mix nodes have both incoming and outgoing edges to other mix nodes and gateways. Gateways serve both as entry and exit points with respect to the mixnet and have incoming and outgoing edges to both mix nodes and clients. For each entity $j$ we define as $P(j)$ the set of its immediate predecessors and $S(j)$ the set of its immediate successors in $G$. The set of forward direction predecessors denoted by $P^{*}(j)$ is defined to be $\emptyset$ for a client $j$, a set of clients in case $j$ is a gateway, and $P^{*}(j)=P(j)\cup_{i\in P^{*}(j)}P^{*}(i)$ for a mix node $j$. A packet route is initiated by a client that is affiliated with a gateway $g_{0}$ and follows a path in $G$ picking outgoing edges and traversing multiple mix nodes. The last mix node in the path is followed by the recipient’s gateway $g_{1}$, who delivers the packet to its intended destination, which could be another client, a service, or a proxy to the open internet.
Client credentials. Subscribed users obtain a bundle of unlinkable credentials (e.g., in the form of a compact e-cash wallet [33] with keys $vk_{\mathsf{long}},sk_{\mathsf{long}}$). Each credential $c$ entitles the client to route up to $S_{c}$ packets through the mixnet.
Network and routing policy. We assume that the mixnet graph $G$ has a layered topology, i.e., $N=LW$ mix nodes are arranged in $L$ layers, with each layer having $W$ nodes – a design that has been shown to offer better anonymity than other topologies [13, 16]. Mix nodes are selected from a pool of at least $N$ eligible nodes and assigned randomly to a layer for a time period (epoch), with the node selection being refreshed every new epoch so that nodes rotate positions and underperforming nodes can be replaced. The set of $W_{G}$ gateways that act as interface between clients and the mixnet is assumed to remain relatively stable over time to minimize the need for clients to switch gateways with the change of epoch, though gateways may be replaced, e.g., if they have low popularity among clients or low reliability. Nodes in the first layer of the mixnet only accept packets from eligible gateways and clients only spend credentials with eligible gateways of their choice.
@@
Blockchain and broadcast channel. We assume that all nodes have access to a blockchain setup that enables them to post transactions that are stored and available long term. Additionally, nodes can use a “layer 2” ephemeral broadcast channel that enables one-to-many information dissemination for information that only needs to be available for a limited time (e.g., a few epochs). Using this layer will be referred to as “broadcasting.” Finally the blockchain can be used to draw randomness that all parties can receive, i.e., offers a randomness beacon.
@@ 2.2 Threat Model
$\bullet$ The adversary may control multiple clients and a limited number of nodes (mix nodes and gateways), such that at least half the predecessors and the successors of each honest node in $G$ are also honest. Malicious nodes can adaptively drop or substitute packets as well as lie when broadcasting output values. Active malicious behavior is assumed to be arbitrary for clients and utility maximizing for gateways: adversarial gateways follow any strategy as long as it does not decrease their utility. We allow any utility function to be used as long as it is monotonically increasing on a number of indicators — defined precisely in Section 3.2.
$\bullet$ “Honest” nodes are divided in a set of reliable nodes that are consistently online and process all received packets and a set of unreliable nodes which experience downtime, throughput limitations, or other faults that result in packet drops. All honest nodes report true output values in every epoch (unless they are down).
$\bullet$ A credential issuer that issues credentials as requested but is not trusted with respect to the privacy of clients.
$\bullet$ A blockchain and ephemeral broadcast channel that cannot be subverted by the adversary, i.e., they satisfy consistency: honest parties have the same view regarding their contents (perhaps subject to a lag between them due to catching up with protocol messages at different times), as well as liveness, i.e., given sufficient time after a transaction is posted either on the ledger or the ephemeral broadcast channel, the transaction is available for all to have access to. Note that in the case of the ephemeral broadcast channel, unreliable honest parties may miss transmitted messages, while in the case of blockchain all honest parties eventually catch up on all posted transactions. Finally the blockchain offers regularly random “beacon” values that are unpredictable by the adversary well in advance.
@@ 2.3 Properties
There are two cryptographic algorithms that typify every mixnet: the packet encoding algorithm that prepares content to be routed via the mixnet and a packet processing algorithm that is executed at every hop in the packet route. In a subscription-based setting, these two algorithms are further equipped with the characteristic that there is a limited amount of bandwidth that a user is authorized to use. Prior to running the packet encoding algorithm, clients obtain a bundle of $N$ credentials; subsequently, they approach a gateway and spend one of the $N$ credentials, which entitles them to submit $S_{c}$ packets to that gateway, who in turn forwards them to the mix nodes. Packets that originate from a particular credential and gateway and are encoded correctly using the packet encoding algorithm are called legitimate packets. Note that a legitimate packet is well-defined given the gateway identity, the credential keys and the random coins used by client and gateway assuming both are running the mixnet encoding algorithm.
Given this feature, for each epoch, we associate the following quantities with a link $e$: ${s}_{e}$ is the number of legitimate packets that were successfully transmitted, i.e., sent and received across the link $e$, and ${d}_{e}$ is the number of legitimate packets dropped in the link. We consider that ${d}_{(c,g)}=0$ in the link $e=(c,g)$ between client $c$ and gateway $g$, as this is a direct live link where clients can retransmit packets until the gateway acknowledges their reception. Thus, ${s}_{(c,g)}$ includes all the packets generated from the client’s credential $c$. At each next step of the path, the quantities ${s}_{e}$ and ${d}_{e}$ of the incoming and outgoing links of a gateway or node $j$ satisfy the following relation.
$\sum_{i\in P(j)}{s}_{(i,j)}=\sum_{k\in S(j)}({s}_{(j,k)}+{d}_{(j,k)})$ (1)
@@
A mixnet with reliability estimation (REst-mixnet) is a mixnet scheme over a graph $G$ where (1) time is divided in epochs; (2) in each epoch a publicly known number of legitimate packets are introduced through each gateway; and (3) for any adversary $\mathcal{A}$ conforming to the threat model of Section 2.2, given an edge $e=(i,j)$, all participants are able to estimate the link’s reliability ${\rho}_{e}$, defined as the fraction of legitimate packets successfully transmitted from $i$ to $j$, relative to all the legitimate packets received by $i$ to be forwarded to $j$; i.e., including both the ${s}_{e}$ packets that were successfully received by $j$ and the ${d}_{e}$ packets that were dropped before being registered by $j$:
@@
Specifically, anyone with access to the broadcast channel can calculate a pair $(\hat{\rho}_{e},\epsilon_{e})$ so that $|\hat{\rho}_{e}-\rho_{e}|\leq\epsilon_{e}$ with confidence level $Z$.
At the same time, we require the following two security properties.
@@
Privacy. Finally, the main objective of any mixnet is to provide unlinkability between input and output packets, assuming an adversary that knows the sender and receiver sets, observes traffic in all the network links (including between clients and gateways) and even controls some of the participants. Thus, the reliability estimation mechanism should not leak any additional information to the adversary that could be useful to deanonymize packets.
3 Reliability Estimation
Our solution relies on indistinguishable measurement packets that are routed through the mixnet with normal client traffic. These measurement packets are later opened and traced along the paths in the mixnet graph. We lay out the description of our protocol in two parts. First, we describe a sampling protocol that generates measurement traffic and enables mix nodes to present evidence regarding their reliability; this allows the reliability estimation algorithm that pulls together the broadcast information to estimate the ratio of successfully transmitted packets in each link and its associated error given a confidence interval; this approach works in a setting where clients and gateways are semi-honest. Second, we introduce VRF-based routing, which is a cryptographic tool that generates unforgeable measurement packets according to the routing policy and limits attacks by malicious clients and gateways.
3.1 Link reliability estimation
We start by describing our link reliability estimation protocol in the simpler setting where clients and gateways are (semi) honest. Then, we show how VRF-based routing can mitigate malicious deviations by clients and gateways.